OpenStudy (anonymous):

x^2+53=0 I need to solve the equation. I need to type an exact answer using radicals as needed. The solution set is????? Need help.

OpenStudy (anonymous):

not possible, unless you are currently doing imaginary numbers?

OpenStudy (anonymous):

is x a real or complex number? If x is real, then solution set is emty set

OpenStudy (anonymous):

I can express in terms of i.

OpenStudy (anonymous):

the x^2 = -53 x^2 = 53i^2 hence x = +/- ixsqrt(53)

OpenStudy (saifoo.khan):

Got the answer?